# Gradient-Free Adversarial Purification with Diffusion Models

### v1

@@ 5.2.1 CIFAR10
5.2.2 ImageNet
Perturbation-based adversarial attack. Tables 4 and 5 demonstrate that the proposed defense method achieves significantly higher performance in both standard accuracy and robust accuracy. Our defense’s standard accuracy notably surpasses previous work, further validating that adversarial super-resolution effectively leverages prior knowledge from the training dataset to achieve better classification accuracy. Adversarial anti-aliasing proves to be particularly effective on the ImageNet dataset, where the filter successfully blurs adversarial perturbations in the detailed pixels of adversarial examples. The performance against PGD-EOT further validates the effectiveness of our proposed defense pipeline.
Unrestricted adversarial attack. We present the defense performance of various methods against the unrestricted adversarial attack AdvDiff and DiffAttack in Table 6 and 7. The results indicate that current defenses are ineffective against the recently proposed unrestricted adversarial attacks. The high standard accuracy can be attributed to the strong generative performance of benign diffusion models. Our defense method is capable of achieving significantly higher robust accuracy compared to previous defenses while preserving the standard accuracy.
Table 6: The standard and robust accuracy against AdvDiff on the ImageNet dataset.
@@ 5.2.2 ImageNet
Table 7: The standard and robust accuracy against DiffAttack on the ImageNet dataset.
@@
5.3 Time efficiency
We evaluate the average time for defending against one adversarial example as shown in Table 8. The results indicate that our proposed method achieves better robust accuracy with significantly lower time costs, as it does not require any gradient calculations over the diffusion model. Notably, our adversarial anti-aliasing can defend against approximately 57% of adversarial examples in just 3e-3 seconds. Furthermore, we can enhance the defense performance of our method by combining it with previous purification methods, with only a minimal tradeoff in time cost.
Table 8: The average time cost of defending one image against PGD ($\ell_{\text{inf}}=4/255$) on the ImageNet dataset.
@@ 5.3 Time efficiency
5.4 Ablation Study
We perform ablation studies to validate the performance of the proposed methods. We evaluate the defense method against AutoAttack ($\ell_{\text{inf}}=8/255$) on the ImageNet dataset by default.
Adversarial Anti-Aliasing. Despite the satisfactory robustness performance of the proposed adversarial anti-aliasing, the choice of filter settings is critical for optimal defense performance. We present the defense performance with different filters in Figure 5. The results indicate a tradeoff between robust accuracy and standard accuracy. Robust accuracy tends to stabilize when using a filter larger than $3\times 3$ in size. Therefore, it is relatively straightforward to identify a suitable filter with a few attempts. Furthermore, the filter settings are generalized across different adversarial attacks within the same dataset, as demonstrated in Tables 4, 5, and 6.
Table 9: The ablation study of proposed methods.
@@ 5.4 Ablation Study
Method Robust Acc(%)
Nie et al. [24] 43.89
+ Ours 69.44
Song et al. [29] 61.53
+ Ours 72.18
Adversarial Super-Resolution. The proposed adversarial super-resolution achieves a similar purification function to previous diffusion-based adversarial purification methods, but without the need for computationally expensive gradient calculations. Table 5(a) demonstrates that our method slightly outperforms traditional adversarial purification when using anti-aliased adversarial examples as input. However, it is crucial to use anti-aliased adversarial examples for optimal performance in adversarial super-resolution, as we do not account for the adversarial gradient during the super-resolution process.
@@
Adversarial Purification. We can enhance diffusion-based adversarial purification methods from previous works by replacing the adversarial input with the adversarial examples after the proposed purification. The processed adversarial examples are more benign and closer to the clean images, thereby enabling better purification performance, as demonstrated in Table 5(b).
6 Conclusion
@@
In this paper, we present an effective and efficient adversarial defense method against both perturbation-based and unrestricted adversarial attacks. The proposed techniques, adversarial anti-aliasing and adversarial super-resolution, effectively eliminate adversarial modifications and recover benign images with minimal computational overhead. Comprehensive experiments on the CIFAR-10 and ImageNet datasets validate that our proposed defense outperforms state-of-the-art defense methods. Our work demonstrates that simple adversarial anti-aliasing can achieve moderate model robustness with almost no additional cost. Furthermore, the proposed super-resolution method can perform adversarial purification without requiring the calculation of the diffusion model’s gradient. We hope our work will serve as a baseline for the further development of adversarial defenses.

### v2

@@ VI-B1 CIFAR-10
[…]
Unrestricted adversarial attack. Unrestricted adversarial examples on the CIFAR-10 dataset are challenging to defend against, as shown in Table V. Our purification method outperforms the previous adversarial purification approach [12] by an average of 10%, validating the effectiveness of our proposed defense. Unrestricted adversarial examples are generated using the same diffusion models employed for purification, with adversarial guidance injected into the sampling process. Since diffusion-based purification relies on a benign generation process that overlooks adversarial guidance, such perturbations are difficult to remove. In contrast, our proposed defense leverages super-resolution diffusion models with a distinct diffusion process. More effectively, the adversarial anti-aliasing module further mitigates adversarial guidance through its blurring-based purification.
VI-B2 ImageNet
TABLE VII: The standard and robust accuracy against AutoAttack ($\ell_{\infty}=8/255$) on the ImageNet dataset.
Method Target Model Standard Acc(%) Robust Acc(%)
Engstrom et al. [47] ResNet50 62.56 31.06
Wong et al. [51] ResNet50 55.62 26.95
Salman et al. [52] ResNet50 64.02 37.89
Bai et al. [53] ResNet50 67.38 35.51
Nie et al. [11] ResNet50 68.22 43.89
Bai et al. [30] ResNet50 70.41 45.59
Chen et al. [13] ResNet50 74.24 43.39
Song et al. [12] ResNet50 66.92 61.53
Ours ResNet50 75.28 $\pm$ 1.06 67.61 $\pm$ 1.95
[…]
TABLE VIII: The defense performance (Robust Accuracy) against AutoAttack ($\ell_{\infty}=8/255$) on the ImageNet dataset against vision transformers.
Method Vit-B(%) Deit-B(%) Swin-B(%) Average(%)
Nie et al. [11] 32.15 28.32 27.48 29.32
Lee et al. [29] 35.92 31.34 30.87 32.71
Wang et al. [32] 42.71 36.50 37.48 38.90
Song et al. [12] 44.30 35.41 38.52 39.41
Ours 68.42 57.84 60.77 62.34
Perturbation-based adversarial attack. ImageNet serves as a more realistic and challenging benchmark for evaluating purification performance, as it consists of 1,000 diverse object categories and high-resolution image data that better reflect real-world scenarios. As reported in Tables VII and VI, our proposed defense achieves a substantial improvement in both standard accuracy and robust accuracy over existing methods. Notably, the standard accuracy of our defense clearly surpasses prior approaches, demonstrating that adversarial super-resolution effectively leverages prior knowledge learned from the training distribution to enhance classification performance under clean conditions. Beyond clean accuracy, the proposed adversarial anti-aliasing mechanism proves particularly effective on the ImageNet dataset. By applying filtering operations, it suppresses adversarial perturbations that are often hidden in the fine-grained pixel structures of high-resolution adversarial examples. This highlights the importance of incorporating frequency-aware defenses when dealing with datasets characterized by detailed textures and complex spatial patterns. In addition, our deblurring-based fine-tuning method further improves robustness against perturbation-based attacks, confirming its general effectiveness across diverse adversarial threat models.
Table VIII further illustrates the strength of our defense when applied to vision transformers. Compared with state-of-the-art diffusion-based purification methods, our approach consistently achieves substantially higher accuracy, underscoring its adaptability to different backbone architectures and its ability to scale effectively with modern large-capacity models.
Finally, the results against adaptive PGD+EOT attacks provide strong evidence of the overall robustness of our pipeline. While such adaptive attacks are theoretically powerful, their execution requires prohibitive computational overhead in terms of both runtime and memory, rendering them impractical for real-time adversarial scenarios.
TABLE IX: The standard and robust accuracy against AdvDiff on the ImageNet dataset.
@@ VI-B2 ImageNet
[…]
TABLE X: The standard and robust accuracy against DiffAttack on the ImageNet dataset.
@@
[…]
Unrestricted adversarial attack. We evaluate the defense performance of various methods against the unrestricted adversarial attacks AdvDiff and DiffAttack, with results reported in Tables IX and X. The findings demonstrate that existing defenses remain largely ineffective against these recently proposed unrestricted adversarial attacks. For AdvDiff (Table IX), the standard accuracy is measured using images generated by the benign sampling process of the latent diffusion model (LDM). The relatively high standard accuracy observed in this setting is primarily due to the strong generative capability of diffusion models, which are trained on the training set of ImageNet dataset and thus inherently produce samples closely aligned with the training distribution. This alignment enables target classifiers to achieve performance comparable to their training accuracy. Importantly, our defense not only preserves this high standard accuracy but also achieves substantially higher robust accuracy than prior defenses. This improvement arises from leveraging a distinct prior different from the LDM, allowing our method to better resist adversarial guidance. In the case of DiffAttack (Table X), adversarial examples are crafted by inverting latents from the ImageNet validation set, making the resulting attacks more similar to perturbation-based adversarial methods. Despite this difference, our defense continues to achieve state-of-the-art performance. The combination of adversarial anti-aliasing and adversarial super-resolution proves highly effective in mitigating adversarial perturbations, thereby delivering superior robust accuracy without sacrificing clean performance. Overall, these results highlight that our proposed defense pipeline not only addresses the limitations of existing diffusion-based purification methods but also establishes strong robustness against the emerging class of unrestricted adversarial attacks.
VI-C Time efficiency
TABLE XI: The average time required to purify a single AutoAttack ($\ell_{\infty}=8/255$) adversarial example on the CIFAR-10 dataset.
Method Defend Method Time Cost(s) Robust Acc(%)
Nie et al. [11] Diffusion 0.6 71.03
Wang et al. [32] Guided Diffusion 2.2 71.18
Lee et al. [29] Diffusion 1.8 70.47
Song et al. [12] Guided Diffusion 2.6 75.45
Ours Adversarial Anti-Aliasing 1e-5 67.85
+ Adversarial Super-Resolution 0.07 82.02
TABLE XII: The average time required to purify a single PGD ($\ell_{\infty}=4/255$) adversarial example on the ImageNet dataset.
@@ VI-C Time efficiency
We evaluate the average runtime required to defend against a single adversarial example, as reported in Tables XI and XII. The results show that our proposed method achieves higher robust accuracy while incurring significantly lower computational costs, primarily because it avoids gradient calculations over the diffusion model. In particular, the adversarial anti-aliasing module can neutralize a large proportion of adversarial examples in under one second, making it highly practical for real-time applications. It is also worth noting that the time efficiency of diffusion-based purification methods is largely determined by the resolution of the input images (i.e., the dataset), which can substantially affect scalability.
Second, we employ diffusion-based super-resolution instead of diffusion-based image generation. It is well-known that the reverse process of diffusion models is time-consuming and the gradient calculation exacerbates this issue. However, we may not require the gradient guidance to maintain the semantic consistency of the label between the purified examples and adversarial examples, given that we already have a reference adversarial example. Since adversarial perturbations are pixel-wise, we opt for a relatively lightweight generation task, namely super-resolution, which also focuses on pixel modification. The diffusion-based super-resolution method used in this paper requires only tens of steps, compared to the hundreds or thousands of diffusion steps needed in previous works. With these two approaches, we significantly enhance the time efficiency of diffusion-based adversarial purification without compromising defense performance.
@@
TABLE XIII: The defense performance against AutoAttack ($\ell_{\infty}=8/255$) on the ImageNet dataset. Time cost on different filters is not evaluated as the computational overhead of adversarial anti-aliasing can be ignored.
Method Target Model Standard Acc(%) Robust Acc(%) Time(s)
Ours (Default Settings, 15 steps) ResNet50 75.28 67.61 1.1
Ours (4 steps) ResNet50 74.81 65.30 0.4
Ours (50 steps) ResNet50 76.20 68.13 4.0
Ours (Default Settings, $3\times 5$ Mean Filter) ResNet50 75.28 67.61 /
Ours ($5\times 5$ Mean Filter) ResNet50 70.67 68.51 /
Ours ($7\times 7$ Mean Filter) ResNet50 68.31 66.81 /
Ours ($3\times 3$ Gaussian Filter) ResNet50 76.21 66.89 /
Ours ($3\times 5$ Gaussian Filter) ResNet50 73.93 66.71 /
Ours ($5\times 5$ Gaussian Filter) ResNet50 71.45 66.44 /
TABLE XIV: The effectiveness of the integration of proposed adversarial purification methods against AutoAttack ($\ell_{\infty}=8/255$) on the ImageNet dataset.
@@
TABLE XV: The ablation study on the selection of the super-resolution diffusion model against AutoAttack ($\ell_{\infty}=8/255$) on the ImageNet dataset.
Method Robust Accuracy(%) Time(s)
ResShift 256 $\times$ 256 67.61 1.1
ResShift 64 $\times$ 64 55.98 1e-1
LDM-SR 128 $\times$ 128 61.35 9.0
SR3 256 $\times$ 256 67.24 25.1
VI-D Ablation Study
We perform ablation studies to validate the performance of the proposed methods. We evaluate the defense method against AutoAttack ($\ell_{\infty}=8/255$) on the ImageNet dataset by default.
Adversarial Anti-Aliasing. Despite the strong robustness of the proposed adversarial anti-aliasing, the choice of filter configuration plays a critical role in achieving optimal defense performance. Figure 6 and Table XIII report results under different filter settings, revealing an inherent trade-off between robust accuracy and standard accuracy. Specifically, robust accuracy tends to stabilize once the filter size exceeds $3\times 3$, making it relatively straightforward to identify an appropriate setting with only a few attempts. Moreover, the effectiveness of a chosen filter generalizes across multiple adversarial attack types within the same dataset, as demonstrated in Tables VII, VI, and IX. Beyond filter size, the type of filter also substantially influences defense performance. As shown in Table XIII, while Gaussian filters are commonly used for image degradation in training models such as ResShift, their effectiveness in adversarial purification is limited compared to our proposed mean non-square filter. More critically, Gaussian filters can be directly neutralized by adaptive attacks, further reducing their robustness. In contrast, our mean non-square filter provides a stronger and more resilient defense, highlighting the importance of carefully designing the filtering strategy for adversarial anti-aliasing.
Adversarial Super-Resolution. The proposed adversarial super-resolution provides a purification mechanism comparable to prior diffusion-based adversarial purification methods, but with the key advantage of avoiding computationally expensive gradient calculations. As shown in Table XIV, our method slightly outperforms DiffPure when performing adversarial anti-aliasing to adversarial examples. It is important to note that anti-aliasing is essential for achieving optimal performance, since the current super-resolution process does not explicitly account for adversarial gradients. To address this limitation, we plan to incorporate adversarial training into the super-resolution model in future work. In addition, we conduct ablation studies with different super-resolution architectures, as reported in Table XV. The results highlight the effectiveness of ResShift, further validating its suitability for adversarial purification. Table XIII shows that increasing the number of diffusion steps improves both standard and robust accuracy. However, this gain comes at the cost of significantly reduced time efficiency due to the longer reverse generation process.
@@ VI-D Ablation Study
Adversarial Purification. We can enhance diffusion-based adversarial purification methods from previous works by replacing the adversarial input with the adversarial examples after the proposed purification. The processed adversarial examples are more benign and closer to the clean images, thereby enabling better purification performance, as shown in Table XIV. The performance gain in adversarial purification further validates the necessity of pre-processing adversarial examples before feeding them into the diffusion purification model. Reducing the magnitude of adversarial perturbations enables more effective purification through the denoising-like reverse generation process. However, applying diffusion purification to adversarial examples multiple times does not yield higher robust accuracy, since residual perturbations are beyond the purification threshold.
Deblurring Fine-Tuning. Our deblurring fine-tuning enhances the performance of the proposed adversarial purification across diverse datasets. While the baseline approach is trained on ImageNet, the fine-tuned method improves both standard and robust accuracy, particularly on CIFAR-10 dataset. This further strengthens the usability and scalability of our defense in practical scenarios across different training dataset.
VII Limitation
Despite achieving significantly higher time efficiency and better defense performance than previous diffusion-based adversarial purification methods, our defense approach still has several limitations. One drawback is that the quality of adversarial examples after purification can be relatively low compared to previous purification methods. Therefore, we recommend using the purified images primarily for detecting adversarial inputs. Another limitation is that the robust accuracy against universal adversarial examples (UAEs) is still not on par with perturbation-based adversarial attacks. We aim to address and improve these limitations in future work.
VIII Conclusion
@@
In this paper, we present an effective and efficient adversarial defense method against both perturbation-based and unrestricted adversarial attacks with theoretical support. The proposed techniques, adversarial anti-aliasing and adversarial super-resolution, effectively eliminate adversarial modifications and recover benign images with minimal computational overhead. The proposed deblurring fine-tuning further enhances the usability and scalability of our purification method across diverse datasets. Comprehensive experiments on the CIFAR-10 and ImageNet datasets validate that our proposed defense outperforms state-of-the-art defense methods. Our work demonstrates that simple adversarial anti-aliasing can achieve moderate model robustness with almost no additional cost. Furthermore, the proposed super-resolution method can perform adversarial purification without requiring the calculation of the diffusion model’s gradient. We hope our work will serve as a baseline for the further development of adversarial defenses.